# Geometry-Aware Edge Pooling for Graph Neural Networks

### v1

@@ D.1 Correlation between Magnitude and Spread
As stated in Section 2.2, the magnitude and spread of a metric space are closely related with magnitude giving an upper bound for spread when computed from the same positive definite metric space. Across our experiments on real graph datasets, we further find that this bound in practice can be very tight and magnitude and spread measure very related notions of effective size. More specifically, when computing both magnitude and spread from the diffusion distances detailed in Section 2.3, we observe that magnitude and spread almost coincide for all graphs from the NCI1, ENZYME or IMDB-Multi datasets as illustrated in Figure 5. In fact, magnitude and spread correlate almost perfectly across these three graph datasets (Pearson correlation $r^{2}\geq 0.99$). Further, we confirm that across these examples, magnitude is generally greater or equal to spread by a relatively low multiplicative factor close to 1. We therefore find empirical evidence for the fact that spread offers a valid and highly related alternative to magnitude in practice supporting our theoretical analysis of the relationship between spread and magnitude during pooling (Section B.3) as well as our observations on the similar performance of both MagEdgePool and SpreadEdgePool.
Figure 5: Comparison between magnitude and spread computed from diffusion distances for all graphs in three graph datasets, NCI1, ENZYMES and IMDB-Multi.
@@ D.2 Evaluating Computational Efficiency
As detailed in Section B.2, the computational costs of our algorithm are determined by the costs of computing the edge scores used for pooling. To illustrate how this theoretical discussion translates into practice, we now investigate computational costs empirically in comparison to alternative pooling methods considered throughout this study.
Specifically, we first compare the runtimes of training the models used in Section 4.3 to compare the accuracy of different pooling methods across varying pooling ratios. Figure 6 then reports the mean runtime of training the GNN on one CV-fold for different choices of pooling ratios and pooling methods for the NCI1 dataset using either general convolutional layers or GIN layers. All models are trained as specified in Table 1 on a single GPU with 32GB memory.
@@
Notably, we observe that MagEdgePool and SpreadEdgePool overall perform on par with alternative pooling methods in terms of runtimes. SpreadEdgePool has a consistent advantage over MagEdgePool due to the higher computational efficiency of computing spread rather than magnitude. Notice that for increasing pooling ratios, our algorithm re-computes the edge scores repeatedly, leading to a less pronounced decrease in computational costs than alternative methods. Nevertheless, we conclude that it is generally more efficient to apply SpreadEdgePool than to rely on trainable approaches, such as TopK and SAGPool for this dataset, indicating the computational benefit of non-trainable graph pooling operations. To generalise this runtime comparison to other datasets sizes and experimental setups, we note that our algorithm will scale with dataset size as described in Section B.2.
@@
Figure 6: Runtime comparison for training the GNNs reported in the main results for NCI1 using different pooling layers. The left plot shows the mean time in seconds per run using general convolutional layers. The right plot shows the same result using GIN layers.
@@ D.3 Node Feature Preservation and Expressivity
Graph pooling should not only preserve graph structure, but also preserve relevant node feature information during pooling. That is, pooling is frequently used after initial rounds of message passing and data representations learnt by previous layers should be respected and effectively encoded by the pooling procedure [7, 36]. One way of investigating node feature retention during pooling, is to evaluate how well a graph can be reconstructed from its pooled version. We follow the experimental setup by Grattarola et al. [25] to investigate. In particular, this experiment uses a model architecture, similar to the model proposed in Table 1, where each graph gets pooled to around 50% of nodes after the initial MLP and GNN layer. Then, the pooled graphs are up-scaled again by reversing the node selection step used by each pooling layer. From these unpooled graph representations, a further GNN and post-processing MLP layer are trained and the task is set to output the reconstructed node feature representation. This model is trained on each example graph using Adam to minimize the mean squared error (MSE) between the input and output node features with a learning rate of $0.0005$ and early stopping on the training loss with a patience of 1000 epochs and a tolerance of $10^{-6}$. Each experiment is repeated three times across different random seeds. See Grattarola et al. [25] for further explanations on the model architecture and experimental setup.
@@
For this experiment, we expect SpreadEdgePool to perform comparably well as we specifically designed our pooling algorithm so that features are averaged during pooling. Further restricting the number of times a node can be merged effectively prevents the collapse of entire portions of the graph, which aids reconstruction. SpreadEdgePool pooling thus successfully encodes node information while allowing for a flexible choice of pooling ratio. This is confirmed by the results in Table 4, which highlight that SpreadEdgePool overall performs well at the features reconstruction task, especially for the sensor graph reaching low reconstruction errors. Alternative methods, in particular node drop approaches such as TopK and SAGPool, show notably worse feature preservation during this experiment indicating the benefits of more expressive pooling operations, such as SpreadEdgePool.
@@
Note that the results in Table 4 capture one specific aspect of feature preservation, namely how well the features of specific example graphs can be reconstructed. This experiment does not assess the generalisation capability of pooling layers. Further, the experimental setup assumes that it is relevant to preserve all node features during pooling, which might not be realistic in practice, where the aim of pooling could be to solely encode task-relevant feature representations. Nevertheless, as discussed above, this extended experiment gives evidence to support that our proposed pooling algorithm, SpreadEdgePool, is capable of outputting expressive feature representations and aggregates node features in a faithful manner, which is likely one of the reasons for its high performance in graph classification tasks.
@@
Table 4: Mean and standard deviation of the reconstruction MSE for reconstructing the original node positions from the pooled graph representations for different example graphs and pooling methods. Our proposed algorithm, SpreradEdgePool, does well at faithfully encoding the feature representations.
Ring Sensor barbell community erdosrenyi torus
@@ D.4 Overview Experiment
To create the illustration in Figure 1 and visually compare the outputs of different pooling methods, we follow the experimental setup by Grattarola et al. [25] on understanding structure preservation in graph pooling layers. Specifically, we simulate a ring graph with 64 nodes, a barbell graph with 20 nodes on each side connected by 24 nodes in the middle, and a sensor graph with 64 nodes. All trainable pooling layers were then trained in a self-supervised manner to optimise the following spectral loss between the original graph $G$ and the pooled graph $G^{\prime}$:
$\mathcal{L}(G,G^{\prime})=\sum_{i=0}^{d}F^{\top}_{:,i}LF_{:,i}-F^{\prime\top}_{:,i}L^{\prime}F^{\prime}_{:,i}$ (15)
Expanding on the qualitative comparison between the example graphs in Figure 1, Figure 7 shows quantitative structure preservation measures for all example graphs and pooling methods. Specifically, we summarise the spectral loss by Grattarola et al. [25], the spectral distance between the normalised graph Laplacians, and the magnitude difference between the pooled and original graphs as further detailed in Section 4.2.
Figure 7: Structure preservation measures for the examples in Figure 1. Pooling is repeated for three different random seeds and the annotations report the means and standard deviations of the structure preservation scores.
Figure 7 demonstrates that SpreadEdgePool and MagEdgePool do not only reach low magnitude differences across these three example graphs, they also show comparatively low spectral distances supporting our findings in Section 4.2. Further, we observe that methods that show worse visual preservation of graph structures, such as NMF, TopK, DiffPool, and MinCut, also reach consistently higher magnitude differences and spectral distances supporting our claim that these methods fail to faithfully preserve graph structures during pooling to varying extents.
@@ D.5 Graph Classification
Table 5 further reports extended classification results on additional datasets extending on the results shown in Table 1. We chose not to report on these datasets in the main text because they showed fewer and less notable differences between pooling methods. Note that for the open graph benchmark MolHIV dataset, instead of using stratified cross-validation, we evaluate each model across predefined training, test and validation splits and evaluate their performance across 5 random seeds. Further, we report AUROC as the performance metric for MolHIV because is the suggested evaluation metric for this very imbalanced dataset. All other results are reported as in Table 1 via the mean and standard deviation of the test accuracy across 10-fold stratified cross-validation. In agreement with our main results, we observe that MagEdgePool and SpreadEdgePool constitute high performing general-purpose pooling methods that reach top performance across these extended datasets. In particular, for smaller biological datasets, such as MUTAG, BZR, or COX2, pooling via magnitude or spread notably improves on the GNN that uses no pooling layer, which indicates the beneficial effects of structure-aware pooling for graph learning.
Table 5: Classification performance of different pooling layers across multiple datasets. For each dataset, the best performing model is marked in bold and models that do not perform significantly different from the best performing model are coloured green.
@@
D.6 Preserving Graph Structure

### v2

@@ D.1 Correlation between Magnitude and Spread
As stated in Section˜2.2, the magnitude and spread of a metric space are closely related with magnitude giving an upper bound for spread when computed from the same positive definite metric space. Across our experiments on real graph datasets, we further find that this bound in practice can be very tight and magnitude and spread measure very related notions of effective size. More specifically, when computing both magnitude and spread from the diffusion distances detailed in Section˜2.3, we observe that magnitude and spread almost coincide for all graphs from the NCI1, ENZYME or IMDB-Multi datasets as illustrated in Figure˜S.1. In fact, magnitude and spread correlate almost perfectly across these three graph datasets (Pearson correlation $r^{2}\geq 0.99$). Further, we confirm that across these examples, magnitude is generally greater or equal to spread by a relatively low multiplicative factor close to 1. We therefore find empirical evidence for the fact that spread offers a valid and highly related alternative to magnitude in practice supporting our theoretical analysis of the relationship between spread and magnitude during pooling (Section˜B.3) as well as our observations on the similar performance of MagEdgePool and SpreadEdgePool.
Figure S.1: Comparison between magnitude and spread computed from diffusion distances for all graphs in three graph datasets, NCI1, ENZYMES and IMDB-Multi.
@@ D.2 Evaluating Computational Efficiency
As detailed in Section˜B.2, the computational costs of our algorithm are determined by the costs of computing the edge scores used for pooling. To illustrate how this theoretical discussion translates into practice, we now investigate computational costs empirically in comparison to alternative pooling methods considered throughout this study.
D.2.1 Training Costs
We first compare the runtime (in seconds) and memory usage (in MB per cross-validation run) of our pooling methods (MagEdgePool and SpreadEdgePool) to trainable pooling methods in Table˜S.3. Specifically, we train the GNN architecture specified in Section˜C.5.3 and Section˜4.1 using GIN layers across 200 epochs using 10-fold stratified cross-validation. We compare our edge pooling methods (MagEdgePool, SpreadEdgePool) to trainable pooling methods from torch_geometric [^12] https://pytorch-geometric.readthedocs.io/en/latest/ available under an MIT license [26, 27]. (EdgePool, TopKPool, SAGPool) or from torch-geometric-pool [^13] https://github.com/tgp-team/torch-geometric-pool available under an MIT license. (DiffPool, MinCutPool). We record the mean and standard deviation of the runtimes in seconds in Table˜S.3 and GPU memory usage in MB per cross-validation run in Table˜S.4. We note, in particular, that our proposed edge pooling methods, MagEdgePool and SpreadEdgePool, allow for significantly more efficient GNN training than EdgePool as highlighted in Table˜S.3. Similarly, our methods generally improve on the runtimes for dense pooling methods such as DiffPool and MinCutPool. To generalise this runtime comparison to other datasets sizes and experimental setups, we note that our algorithm will scale with dataset size as described in Section˜B.2.
D.2.2 Pre-training Costs
Having observed that pre-computed edge-pooling speeds up GNN training times, we further investigate the computational costs of this preprocessing step by comparing the runtime and memory costs of our methods against other non-trainable pooling operators (NDP, NMF and Graclus) prior to training. Table˜S.5 and Table˜S.6 report the computational costs of computing the pooling assignment for all graphs in the datasets. We observe that SpreadEdgePool is generally more efficient than MagEdgePool. Beyond exact computations of our pooling methods (described in Section˜C.4), Table˜S.5 and Table˜S.6 also present approximate versions that reduce the cost of distance computations. Specifically, these approximate versions, referred to as MagEdgePool* and SpreadEdgePool*, use the minimum distance to the original nodes to update the (diffusion) distances during edge contraction. This leads to a considerable improvement in runtime and memory usage during pre-training. The results reported in this section highlight one of the limitations of our edge pooling approach, namely, it scales in the number of edges as well as in the number of nodes as detailed in Section˜B.2. Nevertheless, our proposed pooling methods still outperform EdgePool in terms of computational efficiency when considering both pre-processing and training costs.
Furthermore, the pre-training costs of our method, range in a number of seconds, need to be computed only once per dataset and the memory requirements remain below what is required by GNN training. Hence, while the scalability to very large graphs is a limitation, we find that our proposed pooling methods scale sufficiently well to standard graph datasets. In practice, based on the computational complexity (Section˜B.2), we recommend that our pooling method is particularly suitable for small to medium graphs that show a certain degree of sparsity rather than being fully connected. As visualised in Figure˜S.2, graphs with up to a few hundred nodes can feasibly be processed with our method in a matter of seconds. For future work, we believe that there is a strong potential for adapting our methods to scale on large graphs. For instance, edge score calculations could be parallised, sampling heuristics could restrict the edge score computations to a subset of candidate edges, or edge scores could be estimated from local subgraphs to improve the computations.
Table S.3: Training times in seconds compared across pooling methods. The fastest methods are marked in bold.
Method DHFR ENZYMES NCI109 Mutagenicity IMDB-BINARY IMDB-MULTI
MagEdge 39.8 ± 6.0 22.0 ± 0.2 186.0 ± 9.1 180.3 ± 29.0 75.4 ± 1.4 89.6 ± 6.1
SpreadEdge 34.8 ± 1.0 23.0 ± 0.7 181.3 ± 29.4 206.7 ± 20.6 62.7 ± 0.9 99.7 ± 4.3
EdgePool 184.1 ± 4.0 209.3 ± 4.1 807.4 ± 22.2 790.7 ± 22.1 362.3 ± 3.4 392.7 ± 11.9
TopKPool 50.7 ± 0.8 24.1 ± 0.6 226.1 ± 21.3 243.8 ± 15.4 88.2 ± 1.1 110.9 ± 2.2
SAGPool 54.4 ± 1.3 25.7 ± 0.4 250.4 ± 25.7 253.0 ± 20.0 92.6 ± 2.1 131.1 ± 1.3
DiffPool 52.2 ± 6.8 36.4 ± 3.3 242.4 ± 12.5 274.0 ± 21.5 106.7 ± 5.5 74.9 ± 2.0
MinCut 33.8 ± 2.5 28.2 ± 1.7 210.3 ± 34.4 201.0 ± 3.7 93.2 ± 1.6 132.8 ± 3.2
Table S.4: Memory usage in MB compared across pooling methods. The most efficient methods are marked in bold.
Method DHFR ENZYMES NCI109 Mutagenicity IMDB-BINARY IMDB-MULTI
MagEdge 84.0 ± 0.1 91.2 ± 2.5 97.0 ± 1.1 94.8 ± 1.0 283.4 ± 28.0 197.6 ± 17.8
SpreadEdge 84.0 ± 0.1 90.6 ± 2.3 96.8 ± 1.0 95.2 ± 1.0 283.4 ± 28.0 197.4 ± 17.8
EdgePool 125.6 ± 10.4 148.4 ± 15.0 118.6 ± 7.7 118.0 ± 7.4 666.6 ± 106.6 526.2 ± 89.2
TopKPool 89.6 ± 9.0 94.4 ± 4.0 105.0 ± 1.4 105.8 ± 6.5 259.4 ± 31.5 193.8 ± 17.5
SAGPool 104.2 ± 0.6 94.2 ± 3.8 105.2 ± 1.4 107.8 ± 8.1 273.4 ± 30.1 209.0 ± 17.9
DiffPool 90.2 ± 10.2 94.6 ± 2.5 103.8 ± 1.1 299.8 ± 31.9 258.0 ± 40.3 240.8 ± 26.4
MinCut 90.6 ± 10.3 94.2 ± 2.4 103.6 ± 1.0 288.4 ± 38.3 258.0 ± 40.3 196.6 ± 26.0
Table S.5: Pre-training times in seconds compared across non-trainable pooling methods. The fastest method is marked in bold. The fastest approximation of our pooling method is marked in italics.
Method DHFR ENZYMES NCI109 Mutagenicity IMDB-BINARY IMDB-MULTI
MagEdge 61.9 197.4 4765.1 678.8 801.8 480.2
MagEdge* 24.6 21.0 81.3 74.9 77.4 69.9
SpreadEdge 66.7 68.3 316.1 208.8 287.2 243.8
SpreadEdge* 10.9 16.0 52.3 55.6 56.0 66.0
NDP 5.6 4.1 37.1 32.6 5.8 7.0
NMF 7.8 10.8 39.0 32.1 8.7 9.0
Graclus 3.8 2.9 54.5 18.2 4.5 6.0
Table S.6: Pre-training memory usage in MB compared across non-trainable pooling methods. The most efficient method is marked in bold. The most efficient approximation of our pooling method is marked in italics.
Method DHFR ENZYMES NCI109 Mutagenicity IMDB-BINARY IMDB-MULTI
MagEdgePool 76.2 163.8 158.9 179.0 162.1 93.1
MagEdgePool* 83.7 34.8 47.0 72.2 11.5 65.2
SpreadEdgePool 61.9 60.1 177.9 93.1 204.1 75.8
SpreadEdgePool* 33.6 54.1 60.5 72.9 53.2 23.8
NDP 3.0 5.0 7.0 13.4 5.1 6.5
NMF 3.8 5.9 39.9 10.5 4.2 5.9
Graclus 15.8 7.5 17.5 59.5 1.0 1.5
Figure S.2: Runtime and memory costs of computing the pooling assignment for Erdős-Rényi graphs with edge probability $0.005$ for increasing numbers of nodes. Lines show the mean and shaded areas the standard deviation across five repeats. SpreadEdgePool is notably faster than MagEdgePool. Distance approximations (marked with *) further speed up computations.
D.2.3 Runtimes across Pooling Ratios
@@
Further, we compare the runtimes of training the models used in Section˜4.3 to compare the accuracy of different pooling methods across varying pooling ratios. Figure˜S.3 then reports the mean runtime of training the GNN on one CV-fold for different choices of pooling ratios and pooling methods for the NCI1 dataset using either general convolutional layers or GIN layers. All models are trained as specified in Section˜4.1 on a single GPU with 32 GB memory. Notably, we observe that MagEdgePool and SpreadEdgePool overall perform on par with alternative pooling methods in terms of runtimes. SpreadEdgePool has a consistent advantage over MagEdgePool due to the higher computational efficiency of computing spread rather than magnitude. Notice that for increasing pooling ratios, our algorithm re-computes the edge scores repeatedly, leading to a less pronounced decrease in computational costs than alternative methods. Nevertheless, we conclude that it is generally more efficient to apply SpreadEdgePool than to rely on trainable approaches, such as TopK and SAGPool for this dataset, indicating the computational benefit of non-trainable graph pooling operations.
@@
Figure S.3: Runtime comparison for training the GNNs reported in Section˜4.1 for NCI1 using different pooling layers. Plots show the mean time in seconds per run using general convolutional layers (left) or GIN layers (right).
@@ D.3 Node Feature Preservation and Expressivity
Graph pooling should not only preserve graph structure, but also preserve relevant node feature information during pooling. That is, pooling is frequently used after initial rounds of message passing and data representations learnt by previous layers should be respected and effectively encoded by the pooling procedure [32]. One way of investigating node feature retention during pooling, is to evaluate how well a graph can be reconstructed from its pooled version. We follow the experimental setup by Grattarola et al. [32] to investigate. In particular, this experiment uses a model architecture, similar to the model proposed in Section˜4.1, where each graph gets pooled to around 50% of nodes after the initial MLP and GNN layer. Then, the pooled graphs are up-scaled again by reversing the node selection step used by each pooling layer. From these unpooled graph representations, a further GNN and post-processing MLP layer are trained and the task is set to output the reconstructed node feature representation. This model is trained on each example graph using Adam to minimize the mean squared error (MSE) between the input and output node features with a learning rate of $0.0005$ and early stopping on the training loss with a patience of 1000 epochs and a tolerance of $10^{-6}$. Each experiment is repeated three times across different random seeds. See Grattarola et al. [32] for further explanations on the model architecture and experimental setup.
@@
For this experiment, we expect SpreadEdgePool to perform comparably well as we specifically designed our pooling algorithm so that features are averaged during pooling. Further restricting the number of times a node can be merged effectively prevents the collapse of entire portions of the graph, which aids reconstruction. SpreadEdgePool pooling thus successfully encodes node information while allowing for a flexible choice of pooling ratio. This is confirmed by the results in Table˜S.7, which highlight that SpreadEdgePool overall performs well at the features reconstruction task, especially for the sensor graph reaching low reconstruction errors. Alternative methods, in particular node drop approaches such as TopK and SAGPool, show notably worse feature preservation during this experiment indicating the benefits of more expressive pooling operations, such as SpreadEdgePool.
@@
Note that the results in Table˜S.7 capture one specific aspect of feature preservation, namely how well the features of specific example graphs can be reconstructed. This experiment does not assess the generalisation capability of pooling layers. Further, the experimental setup assumes that it is relevant to preserve all node features during pooling, which might not be realistic in practice, where the aim of pooling could be to solely encode task-relevant feature representations. Nevertheless, as discussed above, this extended experiment gives evidence to support that our proposed pooling algorithm, SpreadEdgePool, is capable of outputting expressive feature representations and aggregates node features in a faithful manner, which is likely one of the reasons for its high performance in graph classification tasks.
@@
Table S.7: Mean and standard deviation of the reconstruction MSE for reconstructing the original node positions from the pooled graph representations for different example graphs and pooling methods. Our proposed algorithm, SpreradEdgePool, does well at faithfully encoding the feature representations.
Ring Sensor Barbell Community Erdős–Rényi Torus
@@
Figure S.4: Structure preservation measures for the examples in Figure˜1. Pooling is repeated for three different random seeds and the annotations report the means and standard deviations of the structure preservation scores.
@@ D.4 Overview Experiment
Expanding on the qualitative comparison between the example graphs in Figure˜1, Figure˜S.4 shows quantitative structure preservation measures for all example graphs and pooling methods. Specifically, we summarise the spectral loss by Grattarola et al. [32], the spectral distance between the normalised graph Laplacians, and the magnitude difference between the pooled and original graphs as further detailed in Section˜4.2.
Figure˜S.4 demonstrates that SpreadEdgePool and MagEdgePool do not only reach low magnitude differences across these three example graphs, they also show comparatively low spectral distances supporting our findings in Section˜4.2. Further, we observe that methods that show worse visual preservation of graph structures, such as NMF, TopK, DiffPool, and MinCut, also reach consistently higher magnitude differences and spectral distances supporting our claim that these methods fail to faithfully preserve graph structures during pooling to varying extents.
@@ D.5 Graph Classification
Table˜S.8 further reports extended classification results on additional datasets extending on the results shown in Table˜1. We chose not to report on these datasets in the main text because they showed fewer and less notable differences between pooling methods. Note that for the open graph benchmark MolHIV dataset, instead of using stratified cross-validation, we evaluate each model across predefined training, test and validation splits and evaluate their performance across 5 random seeds. Further, we report AUROC as the performance metric for MolHIV because is the suggested evaluation metric for this very imbalanced dataset. All other results are reported as in Table˜1 via the mean and standard deviation of the test accuracy across 10-fold stratified cross-validation. In agreement with our main results, we observe that MagEdgePool and SpreadEdgePool constitute high performing general-purpose pooling methods that reach top performance across these extended datasets. In particular, for smaller biological datasets, such as MUTAG, BZR, or COX2, pooling via magnitude or spread notably improves on the GNN that uses no pooling layer, which indicates the beneficial effects of structure-aware pooling for graph learning.
In order to further clarify the classification performance comparison between methods we visualise the ranking of each model choice via a critical difference diagram. This visualisation uses the Friedman test, a non-parametric test for the performance difference between multiple classifiers with repeat measurements, followed by the Nemenyi post-hoc test that facilitates the comparison across all classifiers [17]. In our case, we use the Python package scikit-posthocs [50] to apply this test to the mean performance scores across datasets. The critical difference plot then links all methods that are found to not be statistically different by a horizontal bar. The results reported in Figure˜S.5 confirm that our methods are amongst the best-performing group of pooling layers for the datasets reported in Table˜1. Similarly, Figure˜S.6 shows the critical difference diagram and mean ranks across all tasks included in Table˜1 or Table˜S.8. This extended ranking further strengthens our finding that MagEdgePool and SpreadEdgePool reach top performance across classification tasks.
Table S.8: Classification performance of different pooling layers across multiple datasets. For each dataset, the best performing model is marked in bold and models that do not perform significantly different from the best performing model are coloured green.
@@
Figure S.5: Critical difference diagram for the classification results from Table˜1. Each label corresponds to a choice of pooling method and the method’s mean rank across classification tasks. Groups of methods that are found to not be statistically different are linked by horizontal bars.
Figure S.6: Critical difference diagram for the classification results from Table˜1 and Table˜S.8. Each label corresponds to a choice of pooling method and the method’s mean rank across classification tasks. Groups of methods that are found to not be statistically different are linked by horizontal bars.
D.6 Comparison with EdgePool
To extend our evaluation, we compare with the pytorch_geometric implementation of EdgePool [20, 21]. To do so, we implement the GNN architecture specified in Section˜4.1 in PyTorch and repeat our main experiment reported in Table˜1. We find that there is no significant difference in performance between our methods and EdgePool when computed on e.g. IMDB-M, where EdgePool reaches an accuracy of 47.6 ± 3.3, or on IMDB-B, where EdgePool reaches an accuracy of 70.9 ± 4.5. As further reported in Table˜S.9, similar patterns hold across datasets with our proposed pooling methods reaching comparable accuracies to EdgePool across datasets. As a major advantage, our methods have notably lower computational costs during training than EdgePool as reported in Section˜D.2 and permit choosing flexible pooling ratios. Our methods thus make edge-contraction pooling scalable to larger datasets and enable significantly faster GNN training. On IMDB-B for example, EdgePool takes 362.2 seconds and 666.6 MB for 200 epochs during training, but our method SpreadEdgePool only requires 62.7 seconds and 283.4 MB. We thus find that learning feature-based edge scores as done by EdgePool, is not necessary to ensure classification performance, but rather adds a notable computational burden.
Table S.9: Mean and standard deviation of the graph classification accuracy of different pooling methods across datasets.
Method ENZYMES PROTEINS Mutagenicity DHFR IMDB-B IMDB-M NCI1 NCI109
No Pooling 87.3 ± 2.5 73.8 ± 0.8 80.1 ± 1.3 71.4 ± 1.9 69.7 ± 0.7 46.0 ± 0.7 76.5 ± 1.8 74.3 ± 2.0
MagEdge 91.5 ± 3.2 76.4 ± 3.9 77.5 ± 2.7 88.0 ± 3.8 72.4 ± 1.7 47.4 ± 1.7 72.7 ± 2.4 73.0 ± 3.3
SpreadEdge 92.8 ± 1.6 75.1 ± 3.1 76.0 ± 4.0 90.7 ± 3.8 71.8 ± 1.5 47.3 ± 1.7 73.4 ± 2.5 71.8 ± 1.8
EdgePool 92.1 ± 1.4 75.8 ± 4.6 75.3 ± 3.1 83.2 ± 1.0 70.9 ± 4.5 47.6 ± 3.3 71.9 ± 3.0 72.7 ± 2.6
Random 88.1 ± 2.2 73.7 ± 1.1 73.1 ± 2.1 82.1 ± 3.3 70.1 ± 0.7 45.6 ± 0.8 71.1 ± 2.2 68.0 ± 3.0
D.7 Comparison with Randomised Pooling
Research by Mesquita et al. [44] has demonstrated that randomised ablations of Graclus and DiffPool using random cluster assignment can reach competitive performance to established pooling methods on frequently used graph classification datasets, such as IMDB-B, PROTEINS, NCI109, DD and MOLHIV. Questioning standard assumptions of hierarchical pooling methods, Mesquita et al. [44] thus question the utility of localised graph pooling. We find that this discussion and the observations by Mesquita et al. [44] are closely connected to the following challenges:
  • Graph learning tasks do not necessarily need the graph structure to reach high performance [15, 7].
  • Evaluating the performance of pooling layers is dependent on the specific GNN architecture and benchmarking practices used [24, 44].
  • Randomised modifications of expressive pooling operators can retain the expressivity of the underlying pooling operation [8].
Hence, if all task-relevant information is already captured by the learnt features before pooling, randomised baselines can reach competitive performance. Nevertheless, destroying the structure of graphs during pooling can reduce the effectiveness of preceding message-passing layers [8]. Our pooling approach aims to mitigate this risk by presenting geometry-aware and expressive edge-pooling methods.
As an ablation study, we add randomised edge pooling to our main experiment (cf. Table˜1) by randomly merging pairs of nodes into super-nodes to pre-compute the pooling assignment and averaging their features during training. We find that for NCI109 randomised pooling reaches an accuracy of 68.0 ± 3.0 as compared to 71.8 ± 1.8 for SpreadEdgePool. On NCI1 random pooling reaches an accuracy of 71.1 ± 2.2 compared to 73.4 ± 2.5 for SpreadEdgePool. Thus, the random baseline performs worse than any alternative pooling method on NCI109 and decreases the accuracy on NCI1 slightly. We hypothesise that this decrease is due to graph structure being relevant for these learning tasks [15] and our pooling method better preserves structural properties. In contrast, we observe fewer performance differences for other standard datasets, such as ENZYMES, PROTEINS, IMDB-B, or IMDB-M. This confirms the findings by Coupette et al. [15] that on these datasets GNNs can reach competitive accuracies even when using completely randomised adjacencies as inputs. We thus find that an expressive aggregation of the node features during pooling can suffice [8] and preserving graph structure might not be necessary for solving these tasks to begin with.
D.8 Preserving Graph Structure